# Large Language Models Miss the Multi-Agent Mark

### v2

@@ 1 Introduction
However, margin: 1 margin: MAS LLMs:
a problematic
definition
labelling MAS LLMs as “Multi-Agent Systems” has already raised concerns in the scientific community (chen2024llmcallsneedscaling, ). Influential frameworks employed in MAS LLMs applications, such as ReAct (yao2023reactsynergizingreasoningacting, ), which in turn leverages methods like Chain of Thought (wei2023chainofthoughtpromptingelicitsreasoning, ), Tree of Thought (yao2023treethoughtsdeliberateproblem, ), etc., are single-agent prompting techniques that overlook concurrency and shared states; [^1] https://gist.github.com/yoavg/9142e5d974ab916462e8ec080407365b At the same time, agentic frameworks developed by large companies are monolithic orchestrators that leverage (and only cite) machine learning research, taking little notice of decades of MAS research.2
@@
We articulate our positionmargin: 1 margin: open issues
in MAS LLMs
by identifying three core aspects of MAS that most MAS LLMs in the literature overlook or violate, as well as an aspect related to benchmarking those systems. We criticise the notions of agents’ social intelligence and environment as proposed in the MAS LLMs literature (Sections 2 and 3), and discuss what is missing in terms of coordination and communication (Section 4). Further, we observe that the interest in open-ended environments and emergent behaviours is not supported by benchmarks to define, identify or measure such emergence; the results are primarily descriptive and risk to over-inflate arguments for LLMs’ general and super-intelligence. (Section 5).
@@
I. Social intelligent agents: LLM agents lack native social behaviour. MAS agents populatemargin: 1 margin: the three aspects of intelligence an environment and receive high-level goals to fulfil. Such goals necessitate the realisation, specification, and completion of other possibly unanticipated sub-tasks. In this context, a high-level goal tests an agent’s reactivity, proactiveness, and social abilities (wooldridge1995intelligentagents, ). A reactive agent dynamically perceives the environment and takes the initiative to satisfy its design objectives. In doing so, a social agent is capable of interacting and intelligently competing with other agents.
@@
In the current MAS LLM literature,margin: 1 margin: the social aspects of intelligence we highlight the lack of consideration given to the social aspects of reactivity and proactiveness. While collaborating and competing are prerequisites of any MAS, most LLMs are fine-tuned as single agents, with no proper multi-agent pre-training procedure (kim2023multipromptercooperativepromptoptimization, ). [^2] https://www.kaggle.com/whitepaper-agents, https://www.anthropic.com/engineering/building-effective-agents, https://openai.com/index/new-tools-for-building-agents/ In other words, LLMs are trained in isolation to respond to users’ requests, rather than to interact with each other. This leads to poor performance and unexpected failures when LLMs are benchmarked to identify other agents’ beliefs, desires and intentions, i.e., with Theory of Mind problems (shapira2023cleverhansneuraltheory, ; ullman2023largelanguagemodelsfail, ).
@@
II. Environment design: MAS LLMs environments are LLM-centric. MAS traditionally modelmargin: 1 margin: MAS or
MAS LLMs
design?
the environment with no strong assumption on the architecture or configuration of the agents that populate it (cossentino2005requirements, ; park2005designing, ; wooldridge2000gaia, ). Conversely, MAS LLMs subvert this approach: the environment design assumes the agents are LLMs that communicate and coordinate via natural language. This Section warns that this assumption overlooks the intrinsic limitations of LLMs and argues for addressing such problems by designing MAS environments that are not LLM-centric.
Consider the environment predictability:margin: 1 margin: non-determinism for an environment that is meant to be deterministic, LLMs are inherently not (lamalfa2023jair, ; Ouyang_2025, ); thus, a fully deterministic environment cannot exist when involving LLMs, providing little control over procedures one wants to guarantee to be safe or terminate in a specific state.
Further,margin: 1 margin: inferring beliefs,
desires and
intentions
agents should receive and consistently maintain a unique representation of the environment (Durfee2001, ; rao1995bat, ), particularly when settings are dynamic and partially observable (10.5555/2967142, ). To foster effective cooperation and competition, that should not reduce to the environment but include representing other agents’ actions, beliefs and intentions. Unfortunately, LLMs are not only known to struggle with inferring beliefs and intentions (sclar2023mindinglanguagemodelslack, ; ullman2023largelanguagemodelsfail, ), a long-standing challenge in the MAS literature (10.1007/978-3-031-45368-7_24, ; shi2025mumatommultimodalmultiagenttheory, ), but also with maintaining a consistent representation of the environment due to issues like hallucinations and memory persistency.
Another critical aspect of MAS is how the environment is representedmargin: 1 margin: environment,
perception and
hallucinations
and then perceived by the agents. In most MAS LLMs settings, the environment is textual or translated as such, with text being the medium of reference for the largest majority of open- and closed-source LLMs. Storing such representations may rapidly exceed a model’s context length and cause hallucinations (li2023camel, ; li2023metaagentssimulatinginteractionshuman, ).
@@
A prototypical scenario in a concurrent MASmargin: 1 margin: concurrency
in MAS
is an agent that processes data generated, at random intervals, by another agent. An error may occur when the first agent receives new data before it has finished processing a previous batch. Standard solutions include storing the data or a mechanism to inform the other system to send the data later.
Asynchronicity is typically absent in MAS LLMs,margin: 1 margin: concurrency
in MAS LLMs
as LLM-agents often operate in strictly sequential pipelines or parallel, rather than as independent, concurrently operating agents. While some works are moving in this direction (ginart2024asynchronoustoolusagerealtime, ), we argue for leveraging the body of knowledge from the field of multi-agent concurrent systems that is operative since the 1990s (see, for example, (135681, ; 10.1007/3-540-58855-8_1, )).
Another fundamental aspect of MAS and MAS LLMs is communication between agents. margin: 1 margin: communication
in MAS LLMs
Most MAS LLMs literature assumes agents communicate via natural language (qin2024largelanguagemodelsmeet, ). This smoothness is a simplification that overlooks the complexity of real agent interaction and decades of research in MAS communication systems and protocols (berna2004communication, ; jang2005efficient, ; saunders1996evolution, ).
Natural language communication is, in fact, inefficient and ambiguous:margin: 1 margin: natural language
is inefficient
in contrast, MAS communication protocols in the form of structured languages (e.g., KQML (kqml1994, ) or Agent-Oriented Programming (shoham1993agent, )) are consolidated performative standards that describe agents’ beliefs, commitments, and actions. In line with the MAS literature, we argue mechanisms to negotiate and implement communication methods that integrate the principles of speech acts (austin1962how, ; shoambook, ) and Gricean maxims (Grice1975-GRILAC-6, ) to minimise the cost of communication and maximise its effectiveness.
@@
IV. MAS LLMs do not quantify the emergence of complex behaviours. Several recent worksmargin: 1 margin: emergent
behaviours
considered MAS LLMs as open-ended systems (al2024projectsidmanyagentsimulations, ; park2023generative, ; xie2024largelanguagemodelagents, ), i.e., scenarios where the long-term evolution of the system produces complex behaviours and solutions (openendedlearningteam2021openendedlearningleadsgenerally, ; wang2019pairedopenendedtrailblazerpoet, ). Intuitively, large-scale, long-term interactions would play a similar role as size (the so-called “scaling laws for LLMs” (kaplan2020scalinglawsneurallanguage, )) in shaping complex behaviour and dynamics that would not emerge otherwise.
Currently,margin: 1 margin: grounding
emergent
behaviours
despite growing hype around the potential of LLMs in MAS contexts, which often exceed their demonstrated abilities (shanahan2023talkinglargelanguagemodels, ), we foresee the risk of a birth and death of interest in emergent behaviours in LLMs. Emergence alone, primarily when arising from loosely constrained prompts and undefined interaction dynamics, further compromised by hallucinations and memory issues in LLMs, does not justify claims of MAS-level coordination or long-horizon planning. These factors make distinguishing between genuine coordination and coincidental or spurious outputs increasingly difficult.
@@ 2 Social Intelligent Agents: LLM Agents Lack Native Social Behaviour
an agent is commonly defined as intelligent if it is reactive, proactive, and shows social behaviours (russel2010, ; wooldridge, ). While LLMs are reactive and proactive, this section discusses why they often fail to be social agents and where this discrepancy originates.
@@
For any agent,margin: 1 margin: reactive LLMs the prerequisite of reactiveness is perception, i.e., interpreting an input to take an associated action. LLMs are reactive agents, i.e., respond to stimuli from a changing environment, with the input being directly injected into the prompt (shankar2024buildingreactive, ; sinha2024realtimeanomalydetectionreactive, ; yang2024text2reaction, ) or retrieved via external information (jin2024reasoninggraspingmultimodallarge, ; sun2024interactiveplanning, ).
@@
The proactiveness of an LLMmargin: 1 margin: proactive LLMs is its capacity to initiate tasks with limited or without human intervention (lu2024proactiveagentshiftingllm, ). Independence eventually arises in LLMs only when the model can self-generate the prompt or self-inject information (ouyang2022traininglanguagemodelsfollow, ). Recent works have focused on proactive LLMs that ask clarification to the user on the instructions provided, retrieve additional information from external resources or are fine-tuned to anticipate the next actions (zeighami2025llmpoweredproactivedatasystems, ; zhang2024askbeforeplanproactivelanguageagents, ).
@@
The third intelligent characteristicmargin: 1 margin: social agents of an agent is the capacity to behave socially, thus to compete or cooperate (conitzer2023foundationsof, ; wooldridge1995intelligentagents, ). Competition and cooperation arise from the capacity to be both reactive and proactive, i.e., react to other agents’ actions and initiate new ones. Crucially, agents must be socially reactive and proactive, i.e., able to grasp and reason about other agents’ goals, negotiate with them, and even enlist their cooperation when needed.
The literature on LLMs includes a substantial bodymargin: 1 margin: lack of social
behaviour and
ToM
@@
of works aimed at developing competitive and cooperative systems of LLMs (liu2023bolaabenchmarkingorchestratingllmaugmented, ). In most of them, the agents’ roles are scripted through prompting (du2023improvingfactualityreasoninglanguage, ; mosquera2024llmaugmentedautonomousagentscooperate, ) or fine-tuned (ma2025coevolvingyoufinetuningllm, ; subramaniam2025multiagentfinetuningselfimprovement, ; yang2024multiobjectivefinetuningenhancedprogram, ), but not natively trained to cooperate or compete with one another.
A recent survey,margin: 1 margin: pre-training or
fine-tuning?
@@
identified that $37\%$ of the failure cases of MAS LLMs are errors caused by inter-agents misalignment or agents’ coordination issues (cemri2025multiagentllmsystemsfail, ) (Figure 2). In the same spirit, (Li_2023, ) shows how multi-agent reinforcement learning agents outperform LLMs at planning tasks, highlighting the role of pre-training. Concurrently, a growing body of research in Machine Theory of Mind (ToM) (rabinowitz2018machinetheorymind, ) evidences how LLMs struggle to express their beliefs, desires, and intentions (d2004dmars, ; rao1995bdi, ), and those of other agents (strachan2024testing, ; ullman2023largelanguagemodelsfail, ; van2023theory, ).
Furthermore, when multiple LLMs interact,margin: 1 margin: LLM agents as
ensembles
their specialisation, achieved through different prompts or fine-tuning, often aims to maximise a single objective such as accuracy, at the expense of other desirable behavioural properties. As a consequence, several MAS LLMs reduce to aggregation mechanisms like majority vote (du2024multiagentsoftwaredevelopmentcrossteam, ; kapoor2024aiagentsmatter, ; li2024agentsneed, ; su2025headsbetteroneimproved, ).
The lack of native social behaviour,margin: 1 margin: native interactive
agents
@@
the tendency to converge to ensembles of LLMs rather than developing concurrent strategies, and the well-known issues of LLMs with ToM, contribute to not making LLMs natively interactive agents; in most agentic tools,2 the environment makes them cooperative (e.g., via an orchestrator, workflows, or by prompting them with each other’s output).
@@ 3 Environment Design: MAS LLMs Environments are LLM-centric
As anticipated in the Introduction,margin: 1 margin: LLM-centric
environments
@@
MAS traditionally model the environment with few assumptions regarding the architectures of the agents that will populate it. On the other hand, MAS LLMs centre around LLM-powered agents, overlooking the interoperability with agents that do not conform to them. Furthermore, the intrinsic limitations of LLMs, such as their proneness to hallucinations (liu2024exploringevaluatinghallucinationsllmpowered, ; zhou2025exploringtheproblem, ), lack of determinism, and long-term memory persistency (Ouyang_2025, ) may hinder the development of the field, in particular in scenarios where safety and time (as measured in seconds wooldridge ) are paramount.
In MAS, an environment is the external context that agents perceive and react to,margin: 1 margin: MAS
environments
@@
and can be characterised across five dimensions (russel2010, ; weyns2004environmentsfor, ). Observability determines the amount of information an agent has about the environment and other agents. Predictability states an environment’s determinism level (from fully deterministic to stochastic). Temporality concerns the dependency between successive environment states. Evolution determines how the environment changes with time (i.e., from discrete or continuous). Manipulation represents the degrees of freedom agents have in interacting and modifying the environment.
Of around $110$ MAS LLMs articlesmargin: 1 margin: MAS LLMs
environments
@@
published between $2023$ and $2025$ and whose results we summarise in Figure 1, most MAS LLMs operate in partially observable, deterministic, temporally dynamic, discrete environments (li2024surveyonllm, ; minaee2025largelanguagemodelssurvey, ) (Further details on the papers in Appendix A). Furthermore, the vast majority of LLMs articles (see Figure 1, bottom-right) employ textual representations. [^4] As communication systems are not intrinsically part of the environment, we further elaborate on the non-sustainability of text as the preferred medium of communication in large networks of MAS LLMs in Section 4.
In a fully observable setting, agents receive the same,margin: 1 margin: observability
in MAS LLMS
@@
shared representations of the environment; conversely, partially observable environments provide partial, unique representations to each agent. In MAS LLMs observability, whether partial or full, comes with some issues due to the nature of LLMs: in fact, these models notoriously fail at inferring other agents’ beliefs, desires, and intentions (sclar2023mindinglanguagemodelslack, ; shapira2023cleverhansneuraltheory, ; ullman2023largelanguagemodelsfail, ). As discussed in Section 2, LLMs tend to integrate all the information they have access to without distinguishing what they know as opposed to what they know what other agents know, the so called problem of $k^{\text{th}}$ order beliefs (premack1978does, ). In this sense, a centralised MAS LLMs where an LLM orchestrates and monitors the others may incur issues regarding attribute who did what or what an agent wants to achieve. Even systems with two LLMs suffer from these issues (li2023camel, ), as we later discuss in the last paragraph of this section.
While most MAS LLMs papers assumemargin: 1 margin: non-determinism
in MAS LLMS
@@
the environment is deterministic and can be modified (i.e., it changes according to the actions the LLMs make), the intrinsic non-determinism of LLMs flaws this setting. In terms of safety, one can design specific procedures, such as safety mechanisms and guardrail measures, that deterministically trigger when particular events happen; on the other hand, non-deterministic LLMs provide no guarantees they will behave accordingly (chen2024agentpoisonredteamingllmagents, ). Even setting the temperature of an LLM to zero or sampling deterministically their strategies is not a solution, as many LLMs are known to behave always non-deterministically (lamalfa2023jair, ; Ouyang_2025, ).
In two popular frameworks,margin: 1 margin: hallucinations in
text environments
@@
Camel (li2023camel, ) and MetaAgent (li2023metaagentssimulatinginteractionshuman, ), the authors illustrate how their models hallucinate and underperform when MAS LLMs are specialised to perform a subtask. In Camel, two LLMs autonomously generate prompts to solve complex tasks, reducing reliance on user input. The authors notice that LLMs can inadvertently swap their roles, generate repeated or non-useful instructions or get stuck in infinite message exchanges. MetaAgent focuses on collaborative agents that accomplish coordination tasks. Through perception, memory, reasoning, and execution modules, LLMs interact with the environment, store valuable information, and learn rewarding skills. The authors show that LLMs deviate from their predefined identities, hallucinating their competence and thus compromising collaboration.

### v3

@@ 1 Introduction
However, labelling MAS LLMs as “Multi-Agent Systems” has already raised concerns in the scientific community (chen2024llmcallsneedscaling, ). Influential frameworks employed in MAS LLMs applications, such as ReAct (yao2023reactsynergizingreasoningacting, ), which in turn leverages methods like Chain of Thought (wei2023chainofthoughtpromptingelicitsreasoning, ), Tree of Thought (yao2023treethoughtsdeliberateproblem, ), etc., are single-agent prompting techniques that overlook concurrency and shared states; [^1] https://gist.github.com/yoavg/9142e5d974ab916462e8ec080407365b At the same time, agentic frameworks developed by large companies are monolithic orchestrators that leverage (and only cite) machine learning research, taking little notice of decades of MAS research.2
@@
We articulate our position by identifying three core aspects of MAS that most MAS LLMs in the literature overlook or violate, as well as an aspect related to benchmarking those systems. We criticise the notions of agents’ social intelligence and environment as proposed in the MAS LLMs literature (Sections 2 and 3), and discuss what is missing in terms of coordination and communication (Section 4). Further, we observe that the interest in open-ended environments and emergent behaviours is not supported by benchmarks to define, identify or measure such emergence; the results are primarily descriptive and risk to over-inflate arguments for LLMs’ general and super-intelligence. (Section 5).
@@
I. Social intelligent agents: LLM agents lack native social behaviour. MAS agents populate an environment and receive high-level goals to fulfil. Such goals necessitate the realisation, specification, and completion of other possibly unanticipated sub-tasks. In this context, a high-level goal tests an agent’s reactivity, proactiveness, and social abilities (wooldridge1995intelligentagents, ). A reactive agent dynamically perceives the environment and takes the initiative to satisfy its design objectives. In doing so, a social agent is capable of interacting and intelligently competing with other agents.
@@
In the current MAS LLM literature, we highlight the lack of consideration given to the social aspects of reactivity and proactiveness. While collaborating and competing are prerequisites of any MAS, most LLMs are fine-tuned as single agents, with no proper multi-agent pre-training procedure (kim2023multipromptercooperativepromptoptimization, ). [^2] https://www.kaggle.com/whitepaper-agents, https://www.anthropic.com/engineering/building-effective-agents, https://openai.com/index/new-tools-for-building-agents/ In other words, LLMs are trained in isolation to respond to users’ requests, rather than to interact with each other. This leads to poor performance and unexpected failures when LLMs are benchmarked to identify other agents’ beliefs, desires and intentions, i.e., with Theory of Mind problems (shapira2023cleverhansneuraltheory, ; ullman2023largelanguagemodelsfail, ).
@@
II. Environment design: MAS LLMs environments are LLM-centric. MAS traditionally model the environment with no strong assumption on the architecture or configuration of the agents that populate it (cossentino2005requirements, ; park2005designing, ; wooldridge2000gaia, ). Conversely, MAS LLMs subvert this approach: the environment design assumes the agents are LLMs that communicate and coordinate via natural language. This Section warns that this assumption overlooks the intrinsic limitations of LLMs and argues for addressing such problems by designing MAS environments that are not LLM-centric.
Consider the environment predictability:; thus, a fully deterministic environment cannot exist when involving LLMs, providing little control over procedures one wants to guarantee to be safe or terminate in a specific state.
Further,. To foster effective cooperation and competition, that should not reduce to the environment but include representing other agents’ actions, beliefs and intentions. Unfortunately, LLMs are not only known to struggle with inferring beliefs and intentions (sclar2023mindinglanguagemodelslack, ; ullman2023largelanguagemodelsfail, ), a long-standing challenge in the MAS literature (10.1007/978-3-031-45368-7_24, ; shi2025mumatommultimodalmultiagenttheory, ), but also with maintaining a consistent representation of the environment due to issues like hallucinations and memory persistency.
Another critical aspect of MAS is how the environment is represented and then perceived by the agents. In most MAS LLMs settings, the environment is textual or translated as such, with text being the medium of reference for the largest majority of open- and closed-source LLMs. Storing such representations may rapidly exceed a model’s context length and cause hallucinations (li2023camel, ; li2023metaagentssimulatinginteractionshuman, ).
@@
A prototypical scenario in a concurrent MAS is an agent that processes data generated, at random intervals, by another agent. An error may occur when the first agent receives new data before it has finished processing a previous batch. Standard solutions include storing the data or a mechanism to inform the other system to send the data later.
Asynchronicity is typically absent in MAS LLMs, as LLM-agents often operate in strictly sequential pipelines or parallel, rather than as independent, concurrently operating agents. While some works are moving in this direction (ginart2024asynchronoustoolusagerealtime, ), we argue for leveraging the body of knowledge from the field of multi-agent concurrent systems that is operative since the 1990s (see, for example, (135681, ; 10.1007/3-540-58855-8_1, )).
Another fundamental aspect of MAS and MAS LLMs is communication between agents. Most MAS LLMs literature assumes agents communicate via natural language (qin2024largelanguagemodelsmeet, ). This smoothness is a simplification that overlooks the complexity of real agent interaction and decades of research in MAS communication systems and protocols (berna2004communication, ; jang2005efficient, ; saunders1996evolution, ).
Natural language communication is, in fact, inefficient and ambiguous: standards that describe agents’ beliefs, commitments, and actions. In line with the MAS literature, we argue mechanisms to negotiate and implement communication methods that integrate the principles of speech acts (austin1962how, ; shoambook, ) and Gricean maxims (Grice1975-GRILAC-6, ) to minimise the cost of communication and maximise its effectiveness.
@@
IV. MAS LLMs do not quantify the emergence of complex behaviours. Several recent works considered MAS LLMs as open-ended systems (al2024projectsidmanyagentsimulations, ; park2023generative, ; xie2024largelanguagemodelagents, ), i.e., scenarios where the long-term evolution of the system produces complex behaviours and solutions (openendedlearningteam2021openendedlearningleadsgenerally, ; wang2019pairedopenendedtrailblazerpoet, ). Intuitively, large-scale, long-term interactions would play a similar role as size (the so-called “scaling laws for LLMs” (kaplan2020scalinglawsneurallanguage, )) in shaping complex behaviour and dynamics that would not emerge otherwise.
Currently, despite growing hype around the potential of LLMs in MAS contexts, which often exceed their demonstrated abilities (shanahan2023talkinglargelanguagemodels, ), we foresee the risk of a birth and death of interest in emergent behaviours in LLMs. Emergence alone, primarily when arising from loosely constrained prompts and undefined interaction dynamics, further compromised by hallucinations and memory issues in LLMs, does not justify claims of MAS-level coordination or long-horizon planning. These factors make distinguishing between genuine coordination and coincidental or spurious outputs increasingly difficult.
@@ 2 Social Intelligent Agents: LLM Agents Lack Native Social Behaviour
In MAS, an agent is commonly defined as intelligent if it is reactive, proactive, and shows social behaviours (russel2010, ; wooldridge, ). While LLMs are reactive and proactive, this section discusses why they often fail to be social agents and where this discrepancy originates.
@@
For any agent, the prerequisite of reactiveness is perception, i.e., interpreting an input to take an associated action. LLMs are reactive agents, i.e., respond to stimuli from a changing environment, with the input being directly injected into the prompt (shankar2024buildingreactive, ; sinha2024realtimeanomalydetectionreactive, ; yang2024text2reaction, ) or retrieved via external information (jin2024reasoninggraspingmultimodallarge, ; sun2024interactiveplanning, ).
@@
The proactiveness of an LLM is its capacity to initiate tasks with limited or without human intervention (lu2024proactiveagentshiftingllm, ). Independence eventually arises in LLMs only when the model can self-generate the prompt or self-inject information (ouyang2022traininglanguagemodelsfollow, ). Recent works have focused on proactive LLMs that ask clarification to the user on the instructions provided, retrieve additional information from external resources or are fine-tuned to anticipate the next actions (zeighami2025llmpoweredproactivedatasystems, ; zhang2024askbeforeplanproactivelanguageagents, ).
@@
The third intelligent characteristic of an agent is the capacity to behave socially, thus to compete or cooperate (conitzer2023foundationsof, ; wooldridge1995intelligentagents, ). Competition and cooperation arise from the capacity to be both reactive and proactive, i.e., react to other agents’ actions and initiate new ones. Crucially, agents must be socially reactive and proactive, i.e., able to grasp and reason about other agents’ goals, negotiate with them, and even enlist their cooperation when needed.
@@
The literature on LLMs includes a substantial body. In most of them, the agents’ roles are scripted through prompting (du2023improvingfactualityreasoninglanguage, ; mosquera2024llmaugmentedautonomousagentscooperate, ) or fine-tuned (ma2025coevolvingyoufinetuningllm, ; subramaniam2025multiagentfinetuningselfimprovement, ; yang2024multiobjectivefinetuningenhancedprogram, ), but not natively trained to cooperate or compete with one another.
@@
A recent survey, identified that $37\%$ of the failure cases of MAS LLMs are errors caused by inter-agents misalignment or agents’ coordination issues (cemri2025multiagentllmsystemsfail, ) (Figure 2). In the same spirit, (Li_2023, ) shows how multi-agent reinforcement learning agents outperform LLMs at planning tasks, highlighting the role of pre-training. Concurrently, a growing body of research in Machine Theory of Mind (ToM) (rabinowitz2018machinetheorymind, ) evidences how LLMs struggle to express their beliefs, desires, and intentions (d2004dmars, ; rao1995bdi, ), and those of other agents (strachan2024testing, ; ullman2023largelanguagemodelsfail, ; van2023theory, ).
Furthermore, when multiple LLMs interact,.
@@
The lack of native social behaviour, the tendency to converge to ensembles of LLMs rather than developing concurrent strategies, and the well-known issues of LLMs with ToM, contribute to not making LLMs natively interactive agents; in most agentic tools,2 the environment makes them cooperative (e.g., via an orchestrator, workflows, or by prompting them with each other’s output).
@@ 3 Environment Design: MAS LLMs Environments are LLM-centric
As anticipated in the Introduction, MAS traditionally model the environment with few assumptions regarding the architectures of the agents that will populate it. On the other hand, MAS LLMs centre around LLM-powered agents, overlooking the interoperability with agents that do not conform to them. Furthermore, the intrinsic limitations of LLMs, such as their proneness to hallucinations (liu2024exploringevaluatinghallucinationsllmpowered, ; zhou2025exploringtheproblem, ), lack of determinism, and long-term memory persistency (Ouyang_2025, ) may hinder the development of the field, in particular in scenarios where safety and time (as measured in seconds wooldridge ) are paramount.
@@
In MAS, an environment is the external context that agents perceive and react to, and can be characterised across five dimensions (russel2010, ; weyns2004environmentsfor, ). Observability determines the amount of information an agent has about the environment and other agents. Predictability states an environment’s determinism level (from fully deterministic to stochastic). Temporality concerns the dependency between successive environment states. Evolution determines how the environment changes with time (i.e., from discrete or continuous). Manipulation represents the degrees of freedom agents have in interacting and modifying the environment.
@@
Of around $110$ MAS LLMs articles). Furthermore, the vast majority of LLMs articles (see Figure 1, bottom-right) employ textual representations. [^4] As communication systems are not intrinsically part of the environment, we further elaborate on the non-sustainability of text as the preferred medium of communication in large networks of MAS LLMs in Section 4.
@@
In a fully observable setting, agents receive the same, shared representations of the environment; conversely, partially observable environments provide partial, unique representations to each agent. In MAS LLMs observability, whether partial or full, comes with some issues due to the nature of LLMs: in fact, these models notoriously fail at inferring other agents’ beliefs, desires, and intentions (sclar2023mindinglanguagemodelslack, ; shapira2023cleverhansneuraltheory, ; ullman2023largelanguagemodelsfail, ). As discussed in Section 2, LLMs tend to integrate all the information they have access to without distinguishing what they know as opposed to what they know what other agents know, the so called problem of $k^{\text{th}}$ order beliefs (premack1978does, ). In this sense, a centralised MAS LLMs where an LLM orchestrates and monitors the others may incur issues regarding attribute who did what or what an agent wants to achieve. Even systems with two LLMs suffer from these issues (li2023camel, ), as we later discuss in the last paragraph of this section.
@@
While most MAS LLMs papers assume the environment is deterministic and can be modified (i.e., it changes according to the actions the LLMs make), the intrinsic non-determinism of LLMs flaws this setting. In terms of safety, one can design specific procedures, such as safety mechanisms and guardrail measures, that deterministically trigger when particular events happen; on the other hand, non-deterministic LLMs provide no guarantees they will behave accordingly (chen2024agentpoisonredteamingllmagents, ). Even setting the temperature of an LLM to zero or sampling deterministically their strategies is not a solution, as many LLMs are known to behave always non-deterministically (lamalfa2023jair, ; Ouyang_2025, ).
@@
In two popular frameworks, Camel (li2023camel, ) and MetaAgent (li2023metaagentssimulatinginteractionshuman, ), the authors illustrate how their models hallucinate and underperform when MAS LLMs are specialised to perform a subtask. In Camel, two LLMs autonomously generate prompts to solve complex tasks, reducing reliance on user input. The authors notice that LLMs can inadvertently swap their roles, generate repeated or non-useful instructions or get stuck in infinite message exchanges. MetaAgent focuses on collaborative agents that accomplish coordination tasks. Through perception, memory, reasoning, and execution modules, LLMs interact with the environment, store valuable information, and learn rewarding skills. The authors show that LLMs deviate from their predefined identities, hallucinating their competence and thus compromising collaboration.